Using data found in Appendix E of your textbook calculate the nonstandard emf for each of the following reactions if the concentration of each of the ions in these reactions is 0.0007 molar and everything else is standard (use 298 K for the temperature, R = 8.314 J/mol-K, and F = 96,485 C/mol):
(a) 2 Li+(aq) + 1 Sn(s) --> 2 Li(s) + 1 Sn2+(aq)
(b) 1 Zn2+(aq) + 1 Ca(s) --> 1 Zn(s) + 1 Ca2+(aq)
(c) 1 Mg2+(aq) + 1 H2(g) --> 1 Mg(s) + 2 H+(aq)
(d) 2 Al3+(aq) + 3 Pb(s) --> 2 Al(s) + 3 Pb2+(aq)
Answer #1

Soln niven that I all species have equal conen = 0.00of M R28:314 Tlnelik faa6,485 c/ mich. a) 2+ cor) + sovicsom -> pus +15*

om . Real (a) 24,437/€q) +3Pb (8) - 32Al(s) + 3 Pb2+ (aq) Ecelp > 5exi te Reos = 9.130 -1.66V = -1.53V Ercado = till Boi Ince